Solve. 3x-7x=0 x^2+4=4x
 Thank YOU
: Solve. 3x-7x=0 x^2+4=4x
 Thank YOU

Answer by jim_thompson5910(9166)

# 1



3x-7x=0 Start with the given equation.


-4x=0 Combine like terms on the left side.


x=(0)/(-4) Divide both sides by -4 to isolate x.


x=0 Reduce.


----------------------------------------------------------------------

Answer:

So the answer is x=0





# 2



x^2+4=4x Start with the given equation.


x^2+4-4x=0 Subtract 4x from both sides.


x^2-4x+4=0 Rearrange the terms.


Notice we have a quadratic equation in the form of ax^2+bx+c where a=1, b=-4, and c=4


Let's use the quadratic formula to solve for x


x = (-b +- sqrt( b^2-4ac ))/(2a) Start with the quadratic formula


x = (-(-4) +- sqrt( (-4)^2-4(1)(4) ))/(2(1)) Plug in a=1, b=-4, and c=4


x = (4 +- sqrt( (-4)^2-4(1)(4) ))/(2(1)) Negate -4 to get 4.


x = (4 +- sqrt( 16-4(1)(4) ))/(2(1)) Square -4 to get 16.


x = (4 +- sqrt( 16-16 ))/(2(1)) Multiply 4(1)(4) to get 16


x = (4 +- sqrt( 0 ))/(2(1)) Subtract 16 from 16 to get 0


x = (4 +- sqrt( 0 ))/(2) Multiply 2 and 1 to get 2.


x = (4 +- 0)/(2) Take the square root of 0 to get 0.


x = (4 + 0)/(2) or x = (4 - 0)/(2) Break up the expression.


x = (4)/(2) or x =  (4)/(2) Combine like terms.


x = 2 or x = 2 Simplify.


So the only answer is x = 2